

********************************************************************
UARS WEEKLY STATUS REPORT - 28 MAY 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data without
apparent problems.
We are evaluating the recently reprocessed V20 products.
***********
PEM
***********
NOTHING RECEIVED.
***********
HALOE
***********
NOTHING RECEIVED.
***********
MLS
***********
NOTHING RECEIVED.
***********
HRDI
***********
NOTHING RECEIVED.
***********
WINDII
***********
NOTHING RECEIVED.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
5 days CLA (7.0) PROD (L1-L3) (22R-17C-4F=1L)
34 days PEM (3.32) PROD (OA-L2) (753R-752C-1F=0L)
5 days MLS (4.22) PROD (L1-L3) (JOB 2) (5R-5C=0L)
2 days SOL (4.62) PROD (L2-L3) (TEST) (2R-2F=0L)
72 days WIN (5.11) PROD (L1-L3) (1540R-584C-4F=952L)
The CDHF completed a PEM 3.32 (OA-EXTRACT), Job 2 special request for
a total of 752 days.
The CDHF began the data migration from the current MSS to the new storage
silo on 05/14/99. Two Level 0 platters have been successfully migrated.
DAAC data created for the month of March, 1999 has been loaded into
the DAAC Transfer Tables for HALOE, NMC, PEM, and UKMO.
SOLSTICE Wall Clock time increase. "Several meetings were held with Barry
Knapp to discuss approaches to solving the SOLSTICE problem. Plan to monitor
Ingres file catalogue. SOLSTICE plans a new release in June and intends to
request a mission reprocessing early this summer when the wall clock issue is
resolved."
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
This report covers 15 May 1999 (Orbit 41934, GMT Day 135) through 21 May
1999 (GMT Day 141).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 137/1320 41972 Off
WINDII 137/2006 41976 On
WINDII 137/2150 41977 Off
PEM 138/1438 41988 AXIS Off
PEM 138/2301 41993 Off
HRDI 139/1003 42000 Low Power
MLS 139/1323 42002 Off
WINDII 139/2122 42007 On
MLS 140/1835 42020 On
MLS 141/0119 42024 Off
HRDI 141/1702 42034 Day Only
The beta angle ranged between the angles of 55.9 degrees to 77.1 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 0 degrees on 11 June 1999 (DOY 162).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.7 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.5 to 0.7 degrees C. Additional battery performance data
for each battery on May 14 1999 and May 21 1999 are provided in Appendix D.
The clock error ranged from 11.5 to -10.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% WINDII 05
% MLS 03
% FOT 02
% PEM 02
% HRDI 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/117 - 2287.496987
Transponder B: 99/118 - 2287.498739
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 12 minutes, 50.324 seconds (an increase of 10.976 seconds since
the last report period). This is a 0.0361 percent data loss which equals a
99.9639 percent data capture for the mission.
The increase of 10.976 seconds occurred over two (2) days:
7 1.824 seconds on Day 135 due to a data hit,
7 9.152 seconds on Day 140 due to no overlap.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The end-to-end testing has been delayed to allow additional time for FDF to
create special Y2K planning aids.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL - No change
MEETINGS
FOT attended the weekly IMOC meetings on 21 May 1999.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 11
June 1999 (DOY 162).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 May 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41968 137/0721 99-170 WINDII WINDII Err Fix
41972 137/1320 99-171 WINDII Power OFF
41976 137/2006 99-172 WINDII Power ON
41977 137/2150 99-173 WINDII Power OFF
41988 138/1438 99-174 OBC Clk Adj 20.6 to 41.2 msec/day
41988 138/1438 99-175 PEM AXIS Power OFF
41993 138/2301 99-176 PEM Power OFF
42000 139/1003 99-177 HRDI Power OFF
42002 139/1323 99-178 MLS Power OFF
42007 139/2122 99-179 WINDII Power ON
42020 140/1835 99-167 MLS Power ON
42024 141/0119 99-180 MLS Power OFF due to s/c low power
42034 141/1702 99-181 HRDI Power ON (day only)
42035 141/1806 99-182 OBC Clk Adj fr 41.2 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 May 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIQBCURRENT Y-LO WINDII Quiet Bus Current 42007-8
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 May 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
41988 138/14:42 (18 May 99) from 20.6 to 41.2 msec/day
42035 141/18:13 (21 May 99) from 41.2 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 May 1999
14 May 1999 - GMT Day 134 Beta = 55.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -78.4/-112.0 -3.68/-4.94 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-28.0 +5.15/+3.87 +24.8/-16.4 23.8 1.013 21.0
3 +22.4/-50.4 +4.51/+2.60 +24.8/-17.2 23.8 1.016 20.8
21 May 1999 - GMT Day 141 Beta = 77.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -89.6/-106.4 -2.11/-3.05 +0.0/ +0.0 16.5 0.000 0.0
2 +11.2/-11.2 +2.92/+1.34 +15.2/-16.0 25.1 1.077 12.8
3 +22.4/-44.8 +2.29/+0.71 +15.2/-16.0 25.1 1.081 12.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 May 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-023 FOT see list May 99 14170 Generic Late Acq
SEVNTFW99/21